# Tomographic redshift dipole: Testing the cosmological principle

### v2

@@ Introduction
Survey systematics, including observational strategies, footprint inhomogeneities, and variations in sky and telescope conditions, could introduce measurement bias. To assess and then adjust for this bias, we developed a series of mock catalogs from eBOSS’s EZmock sample and BOSS’s MultiDark-Patchy mocks. Notably, we utilized the realistic version of EZmock, which incorporates observational systematics, unlike the MultiDark-Patchy mocks that lack this aspect. Therefore, the findings related to the BOSS catalogs might be influenced by observational systematics. Our collection of mock catalogs was created by Doppler boosting 100 randomly selected mocks from both the eBOSS and BOSS surveys in 12 different directions, uniformly distributed across the sky. This process, utilizing a velocity of 370 km/s consistent with the kinematic interpretation of the CMB dipole, produced a total of 1200 mock catalogs for each tracer, summing up to 6000 mock catalogs overall.
Although we detected a relatively minor bias, we proceeded to apply a bias correction technique utilizing bi-linear interpolation of the bias vector. By subtracting the anticipated bias vector from the observed dipole measurement, we derived the de-biased value. This correction method was applied across both individual tracers and their combinations, as detailed in Methods. Fig. 4 demonstrates this process for the combined cases, by contrasting the average biased and de-biased dipole vectors against the benchmark dipole set in the simulations. See Extended Data Fig. 2 for the individual tracers’ cases. The corrected bias for each Cartesian component agrees with the benchmark values. This agreement was verified by comparing the mean difference vectors to the standard errors derived from 100 mocks per direction. Furthermore, we validated the estimator’s efficacy on the original 5000 mocks without Doppler effect, achieving the predicted outcomes after employing the bias correction, as illustrated in Extended Data Fig. 4.
Refer to captiona) Results in galactic coordinatesRefer to captionb) Results in Cartesian coordinates
Figure 5: Final dipole measurements. (a) Mollweide projection displaying the dipole measurements in galactic coordinates, with contours representing the $1\sigma$ and $2\sigma$ confidence regions for combined measurements. The star symbol denotes the position of the CMB dipole for reference. (b) Probability distribution ridgeline plots in Cartesian coordinates, with dashed lines marking the CMB dipole values and dots indicating the $1\sigma$ confidence intervals.
@@
When observed within the same redshift range, different tracers are expected to exhibit a similar clustering dipole, $\boldsymbol{\Delta_{1,\rm{int}}}$, which could lead to correlations between measurements not accounted for by our estimator. Despite our theoretical assessment indicating that the intrinsic dipole’s influence is minor (less than 2% of the anticipated dipole), we utilized the BOSS and eBOSS mock datasets to evaluate this effect through an alternative approach. Our empirical results reaffirm the theoretical prediction, establishing that the clustering dipole’s effect is indeed negligible, as depicted in Extended Data Fig. 3.
@@
To accurately estimate the uncertainty associated with the dipole estimation, we applied a Doppler boost to the 1000 BOSS and eBOSS mock catalogs based on the best-fit model derived from the real dataset (for a total of 5000 mocks). We then processed these boosted mocks through the complete pipeline, effectively sampling the measurement distribution to determine the confidence intervals. Notably, the distributions obtained are consistent with the unbiased best-fit dipoles, serving as a further critical consistency check of the effectiveness of our bias removal technique across various dipole amplitudes, demonstrating that the bias correction is not significantly influenced by the absolute value of the dipole, including those different from the CMB-derived expectation used to create the bias maps.
@@
Finally, given our utilization of a finite set (1000) of mock catalogs for statistical analysis, we ensure the robustness of our findings by employing two complementary methods, both leading to similar conclusions: a non-parametric Kernel Density Estimation (KDE) with a Gaussian kernel and a parametric best-fit distribution analysis using the Distfit Python module40. This approach was consistent across error estimation and the determination of two-tailed significance regarding potential discrepancies with the CMB dipole. Our primary analysis leverages KDE. In agreement with our frequentist methodology for assessing uncertainties using mock catalogs, central values and $1\sigma$ uncertainties are determined using quantiles: the 50th percentile for the central value, and the 16th and 84th percentiles for the lower and upper uncertainty bounds, respectively.
Results and Discussion
@@
[…]
To evaluate the two-tailed significance, we compare the dipoles to the CMB value in Cartesian coordinates using two methods: a non-parametric Kernel Density Estimation (KDE) with a Gaussian kernel and a parametric best-fit distribution analysis conducted using the Distfit Python module40. The table additionally lists the effective redshift for each scenario. For BOSS and eBOSS data, this corresponds to the weighted average redshift, which is equivalent to the monopole.
The analysis pipeline used in this study, including the evaluation of potential tension with the CMB dipole, was exclusively developed and validated using mock catalogs, ensuring that real data remained unexamined until after the methodology was established. Maintaining data blinding is critical to avoid reliance on a posteriori statistics, which might inadvertently skew the findings. In total, we processed 16000 mock catalogs, each subdivided into thousands of finely segmented redshift bins. The computational demand of this extensive study amounted to approximately 200000 CPU hours. For a visual representation of the data analysis pipeline, please refer to Extended Data Fig. 5.
Our findings are depicted in Fig. 5 and detailed in Table 1. The confidence contours illustrated in Fig. 5a were generated with the ligo.skymap Python module [^1] See documentation at lscsoft.docs.ligo.org/ligo.skymap., a tool frequently utilized in the analysis of gravitational wave observations. Fig. 5b showcases the probability distributions ascertained through KDE. Table 1 lists the confidence intervals for both the magnitude and galactic coordinates of the velocities determined using the various catalogs and their combinations, alongside the significance of discrepancies relative to the CMB dipole. These significances were calculated for each Cartesian component and combined employing Fisher’s method41, see Methods.
Redshift dipole measurements indicate agreement with the CMB dipole, exhibiting a tension of $0.8\sigma$ for eBOSS data ($2.20.6<z<$ 2.2), $2.1\sigma$ for BOSS data ($0.60.2<z<$ 0.6), and $1.4\sigma$ for combined eBOSS+BOSS data ($2.20.2<z<$ 2.2). The most notable, albeit moderate, tension with the anticipated dipole arises from BOSS data within the redshift range $0.60.2<z<$ 0.6, where our findings suggest a deviation towards the southern hemisphere, despite the dipole’s magnitude closely matching expectations. This discrepancy slightly amplifies the overall tension between the combined measurement and the CMB dipole, while also decreasing the dipole magnitude due to the lower latitude of BOSS’s measured dipole. As previously highlighted, the BOSS mock catalogs did not incorporate realistic systematic effects, rendering the bias removal method unable to account for these unmodeled systematics, though they do incorporate veto weights to flag areas potentially affected by systematics. This consideration is relevant when evaluating the combined results’ overall increased tension and reduced dipole magnitude. However, it’s important to highlight that the bias correction applied across all scenarios resulted in only minor adjustments to the observed dipoles, revealing no significant evidence of systematic effects within the eBOSS or BOSS datasets.
[…]
Our analysis reveals that our motion with respect to distant galaxies and quasars is in close agreement with that suggested by the CMB dipole, as illustrated in Fig. 5 in both Galactic and Cartesian coordinates. This concordance implies that matter at cosmological distances of 1–5 Gpc shares the same rest frame as the CMB, located approximately 14 Gpc away, offering substantial support to the cosmological principle, which advocates for large-scale homogeneity and isotropy. However, significant tensions highlighted by previous studies suggest that the possibility of deviations from the cosmological principle cannot be fully excluded by our measurements alone. Progress in this area will require careful consideration of potential overlooked systematic factors in analyses based on number counts and redshifts, such as redshift evolution modeling, biases in estimators, or underestimation of uncertainties42, 43, 44, 45, 46. Furthermore, validating the analysis pipeline using realistic mock catalogs and conducting blind data analyses, as undertaken in this study, remains critical.
Cosmology aims to comprehend the universe at its broadest spatial and temporal scales, relying on foundational principles to simplify our understanding of spacetime. Central to modern cosmology is the assumption of the universe’s statistical homogeneity and isotropy at large scales. The assumption’s validity stems from the observed near-isotropy of the universe and the Copernican principle, positing no special vantage point for observers: if all observers perceive the universe as isotropic, it must also be homogeneous. Our study tackles, therefore, a fundamental question on our place in the universe, reinforcing the mathematical and philosophical foundations of our standard cosmological model.
Methods
Estimator
We employ a least squares estimator to measure the dipole $\boldsymbol{\Delta}_{1}$ for each bin, based on the peculiar velocity hypothesis $\boldsymbol{\Delta}_{1}=-\boldsymbol{\beta}$. This calculation involves comparing the observed redshift distribution to the expected one for a theoretically isotropic $z$-field, boosted as per Eq. (1):
$\chi^{2}_{\rm bin}(\boldsymbol{\Delta}_{1})=\frac{\sum_{i}^{N}w_{i}^{2}\left[z_{i}-\Delta^{dD}_{0,{\rm bin}}\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})\right]^{2}}{\sum_{i}^{N}w_{i}^{2}},$ (6)
$w_{i,\mathrm{eBOSS}}$ $=w_{i,\mathrm{sys}}\cdot w_{i,\mathrm{noz}}\cdot w_{i,\mathrm{fc}}\,,$ (7)
$w_{i,\mathrm{BOSS}}$ $=w_{i,\mathrm{sys}}\cdot(w_{i,\mathrm{noz}}+w_{i,\mathrm{fc}}-1)\,,$ (8)
accounting for redshift failures ($w_{i,\mathrm{noz}}$), fiber collisions ($w_{i,\mathrm{fc}}$), and imaging systematics ($w_{i,\mathrm{sys}}$). Unlike other estimators, we consider the modulation around the “de-Dopplered” monopole $\Delta^{dD}_{0,{\rm bin}}$ of the respective bin, which is the average $z$ with the Doppler modulation removed:
$\Delta^{dD}_{0,{\rm bin}}=\sum_{i}^{N}w_{i}z_{i}\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})^{-1}\bigg{/}\sum_{i}^{N}w_{i}\,.$ (9)
Spectroscopic redshift measurement errors are negligible47 and thus not considered. As discussed in Data, it’s necessary to analyze objects in the Northern and Southern Galactic Caps (NGC and SGC) separately due to distinct monopoles. Therefore, to find the complete sample’s best-fit dipole $\boldsymbol{\Delta}_{1}$, we minimize across all bins according to:
$\chi^{2}(\boldsymbol{\Delta}_{\rm 1})=\sum_{\rm bin}\chi^{2}_{\rm bin,NGC}(\boldsymbol{\Delta}_{1})+\sum_{\rm bin}\chi^{2}_{\rm bin,SGC}(\boldsymbol{\Delta}_{1})\,.$ (10)
We also note that in our analysis, we disregard terms of the order $\mathcal{O}(\beta^{2})$.
Data combination
To combine the results relative to the individual tracers and obtain the eBOSS, BOSS and eBOSS+BOSS results, we implemented the following steps: i) We determined the probability distributions $\{P^{\rm case}_{x},P^{\rm case}_{y},P^{\rm case}_{z}\}$ for each Cartesian component of the dipole vector $\{\Delta^{\rm case}_{1,x},\Delta^{\rm case}_{1,y},\Delta^{\rm case}_{1,z}\}$, across all individual cases; ii) We then multiplied the distributions of the Cartesian components together, followed by renormalization. For instance, prior to renormalization,
$\!\!\!\!\!P^{\text{\tiny{BOSS}}}_{x}\!$ $=\!P^{\text{\tiny{CMASS BOSS}}}_{x}\!\times\!P^{\text{\tiny{LOWZ BOSS}}}_{x}\,$
$P^{\text{\tiny{eBOSS}}}_{x}\!$ $=\!P^{\text{\tiny{QSO eBOSS}}}_{x}\!\times\!P^{\text{\tiny{LRG eBOSS}}}_{x}\!\times\!P^{\text{\tiny{CMASS eBOSS}}}_{x}$
$P^{\text{\tiny{eBOSS+BOSS}}}_{x}\!$ $=\!P^{\text{\tiny{eBOSS}}}_{x}\!\times\!P^{\text{\tiny{BOSS}}}_{x}\,,$ (11)
Fisher’s method
Fisher’s method is utilized to consolidate results from multiple independent tests that assess the same hypothesis41. It aggregates the $p$-values from each test into a single test statistic, $X^{2}_{2k}$, through the equation:
$X^{2}_{2k}=-2\sum_{i=1}^{k}\ln p_{i},$ (12)
where $p_{i}$ represents the $p$-value for the $i$-th hypothesis test. Assuming the null hypothesis is valid and the $p_{i}$ values (or their corresponding test statistics) are independent, $X^{2}_{2k}$ follows a chi-squared distribution with $2k$ degrees of freedom, $k$ being the count of tests being combined. This can be used to determine the combined $p$-value, which we express in $\sigma$ units.
Correlations
Here, we used the BOSS and eBOSS mock datasets to assess whether different tracers display a comparable clustering dipole, attributable to their sampling of the same underlying dark matter density and velocity field. This effect potentially induces correlations between measurements. EZmock and MultiDark-Patchy individual tracers’ mocks, sharing identical ID numbers, originate from the same dark matter realization, ensuring they exhibit a similar clustering dipole within the same redshift range. We evaluated the dipole for each simulation, confirming that the measurements agree with an expected null dipole, as depicted in Extended Data Fig. 4. Based on these results, we calculated the correlation between different catalogs, finding that, as illustrated in Extended Data Fig. 3, the correlation is insignificant. This implies that the primary source of uncertainty is attributed to the finite tracers’ number density.
Bias removal
To construct the bias map, we Doppler boosted 100 mocks randomly chosen from both the eBOSS and BOSS surveys across 12 distinct directions, located at the pixel centers in the HEALPix $N_{\rm side}=1$ scheme, as illustrated in Fig. 4. HEALPix48, 49, or Hierarchical Equal Area isoLatitude Pixelization, serves as an efficient system for handling spherical data, extensively employed in astrophysics and cosmology for pixelating the celestial sphere to streamline the analysis and visualization of astronomical datasets. This step assumes a velocity of 370 km/s, in line with the kinematic interpretation of the CMB dipole, yet, as noted in the main text, the bias map remains reliable across various dipole magnitudes, including the null case.
To compute the bias vector, we first estimate the dipole for each of the 100 simulations per direction (1200 simulations per tracer). We then derive the average difference vectors by subtracting the fiducial dipole vector $\boldsymbol{\Delta}^{\rm{fid,i}}_{1}$ for direction $i$ from the simulations’ average dipole vector $\langle\boldsymbol{\Delta}^{\rm{sim,i}}_{1}\rangle$. These average difference vectors, through their Cartesian components, establish the pixel values of the three bias maps for each tracer. Subsequent bi-linear interpolation of these bias maps for all Cartesian components is conducted using the get_interp_val function from the healpy Python module49. Through this interpolation, we can de-bias the measured dipole, ensuring accurate analysis and interpretation of our findings.
BOSS mocks, while similar to EZmocks, do not incorporate systematic effects. As a result, the weights applied in the estimation of the bias maps are defined as follows:
$w^{\rm{mocks}}_{i,\mathrm{BOSS}}=w_{i,\mathrm{veto}}\cdot w_{i,\mathrm{fc}},$ (13)
where $w_{i,\mathrm{veto}}$ acts as a flag, being set to 1 for objects within the “veto mask” and 0 for those outside. The veto mask is employed to exclude sections of the sky that fail to adhere to specific quality standards, thus mitigating areas heavily influenced by systematics. Such standards may encompass regions lacking adequate observational coverage, areas subject to contamination, or any other conditions potentially compromising data quality.
The described method outlines the process for generating bias maps for individual tracers. To construct bias maps for the aggregate analyses of eBOSS, BOSS, and eBOSS+BOSS, we adopt the strategy used for combining data, as depicted in Eq. (11). Initially, we correct the 100 mock dipole measurements per direction using the bias maps specific to each tracer, then we aggregate these distributions and compare the composite average dipole to the fiducial one. The resultant difference vector forms the basis of the bias map for the combined analysis, which we subsequently interpolate to correct the measured dipoles derived from data combinations.

### v3

@@ 4.2 Bias estimation and removal
4.3 Correlations
@@
When observed within the same redshift range, different tracers are expected to exhibit a similar clustering dipole, $\boldsymbol{\Delta_{1,\rm{int}}}$, which could lead to correlations between measurements not accounted for by our estimator. Despite our theoretical assessment indicating that the intrinsic dipole’s influence is minor (less than 2% of the anticipated dipole), we employed the eBOSS mock datasets for a practical evaluation.
Using these datasets, we assessed whether different tracers within the same redshift range demonstrate a comparable clustering dipole, attributable to sampling the same underlying dark matter density and velocity field. This effect potentially induces correlations between measurements. EZmock individual tracers’ mocks, sharing identical ID numbers, originate from the same dark matter realization, ensuring they exhibit a similar clustering dipole within the same redshift range. We evaluated the dipole for each simulation and calculated the correlation between different catalogs, finding that, as illustrated in Fig. 8, the correlation is insignificant. This implies that the primary source of uncertainty is attributed to the finite tracers’ number density. Our empirical results reaffirm the theoretical prediction, establishing that the clustering dipole’s effect is indeed negligible.
Figure 8: Pearson correlation matrix between $\Delta_{1}$ measurements of different tracers. The correlations were obtained using the eBOSS mocks.
4.4 Error estimation
@@
To accurately estimate the uncertainty associated with the dipole estimation, we applied a Doppler boost to the 1000 BOSS and eBOSS mock catalogs based on the best-fit model derived from the real dataset (for a total of 4000 mocks). We then processed these boosted mocks through the complete pipeline, effectively sampling the measurement distribution to determine the confidence intervals. Notably, the distributions obtained are consistent with the unbiased best-fit dipoles, serving as a further critical consistency check of the effectiveness of our bias removal technique across various dipole amplitudes, demonstrating that the bias correction is not significantly influenced by the absolute value of the dipole, including those different from the CMB-derived expectation used to create the bias maps.
@@
Given our utilization of a finite set (1000) of mock catalogs for statistical analysis, we ensure the robustness of our findings by employing two complementary methods, both leading to similar conclusions: a non-parametric Kernel Density Estimation (KDE) with a Gaussian kernel and a parametric best-fit distribution analysis using the Distfit Python module [44]. This approach was consistent across error estimation and the determination of two-tailed significance regarding potential discrepancies with the CMB dipole. Our primary analysis leverages KDE. In agreement with our frequentist methodology for assessing uncertainties using mock catalogs, central values and $1\sigma$ uncertainties are determined using quantiles: the 50th percentile for the central value, and the 16th and 84th percentiles for the lower and upper uncertainty bounds, respectively.
4.5 Data combination
To combine the results relative to the individual tracers and obtain the eBOSS results, we implemented the following steps: i) We determined the probability distributions $\{P^{\rm case}_{x},P^{\rm case}_{y},P^{\rm case}_{z}\}$ for each Cartesian component of the dipole vector $\{\Delta^{\rm case}_{1,x},\Delta^{\rm case}_{1,y},\Delta^{\rm case}_{1,z}\}$, across all individual cases; ii) We then multiplied the distributions of the Cartesian components together, followed by renormalization. For instance, prior to renormalization,
$P^{\text{\tiny{eBOSS}}}_{x}\!$ $=\!P^{\text{\tiny{QSO eBOSS}}}_{x}\!\times\!P^{\text{\tiny{LRG eBOSS}}}_{x}\!\times\!P^{\text{\tiny{CMASS eBOSS}}}_{x}\;\,,$ (4.7)
The bias removal method of Section 4.2 yields bias maps for individual tracers. To construct bias maps for the aggregate analysis of eBOSS, we adopt the strategy used for combining data, as depicted in Eq. (4.7). Initially, we correct the 100 mock dipole measurements per direction using the bias maps specific to each tracer, then we aggregate these distributions and compare the composite average dipole to the fiducial one. The resultant difference vector forms the basis of the bias map for the combined analysis, which we subsequently interpolate to correct the measured dipoles derived from data combinations.
5 Results and Discussion
The analysis pipeline used in this study, including the evaluation of potential tension with the CMB dipole, was exclusively developed and validated using mock catalogs, ensuring that real data remained unexamined until after the methodology was established. Maintaining data blinding is critical to avoid reliance on a posteriori statistics, which might inadvertently skew the findings. In total, we processed 15000 mock catalogs, each subdivided into thousands of finely segmented redshift bins. The computational demand of this extensive study amounted to approximately 250000 CPU hours. For a visual representation of the data analysis pipeline, please refer to Appendix A.
Figure 9: Final dipole measurements. Mollweide projection displaying the dipole measurements in galactic coordinates, with contours representing the $1\sigma$ and $2\sigma$ confidence regions for combined measurements. The star symbol denotes the position of the CMB dipole for reference.
Our findings are depicted in Fig. 9 and detailed in Table 1. The confidence contours illustrated in Fig. 9 were generated with the ligo.skymap Python module [^2] See documentation at lscsoft.docs.ligo.org/ligo.skymap., a tool frequently utilized in the analysis of gravitational wave observations. Table 1 lists the confidence intervals for both the magnitude and galactic coordinates of the velocities determined using the various catalogs and their combinations, alongside the significance of discrepancies relative to the CMB dipole. These significances were calculated for each Cartesian component and combined employing Fisher’s method [45], see Appendix B.
@@ 5 Results and Discussion
To evaluate the two-tailed significance, we compare the dipoles to the CMB value in Cartesian coordinates using two methods: a non-parametric Kernel Density Estimation (KDE) with a Gaussian kernel and a parametric best-fit distribution analysis conducted using the Distfit Python module [44]. The table additionally lists the effective redshift for each scenario. For BOSS and eBOSS data, this corresponds to the weighted average redshift.
Redshift dipole measurements from the combined eBOSS data ($2.20.6<z<$ 2.2) demonstrate a $2\sigma$ tension with the CMB dipole, indicating broad agreement. Dipole measurements across individual eBOSS catalogs are also mutually consistent. Figure 1 shows a very good agreement ($\sim 1\sigma$ level) between these results and the ones obtained in refs. 21, 17, 13, 18. Our analysis using eBOSS data reveals that our motion with respect to distant galaxies and quasars is in close agreement with that suggested by the CMB dipole. This concordance implies that matter at cosmological distances of 2–5 Gpc shares the same rest frame as the CMB, located approximately 14 Gpc away. Such findings reinforce the mathematical and philosophical foundations of our standard cosmological model, which posits large-scale homogeneity and isotropy.
However, BOSS data ($0.60.4<z<$ 0.6) suggest a deviation towards the southern hemisphere, in strong tension ($5\sigma$) with the CMB dipole. As previously highlighted, the BOSS mock catalogs did not incorporate realistic systematic effects, rendering the bias removal method unable to account for these unmodeled systematics, though they do incorporate veto weights to flag areas potentially affected by systematics. While our bias correction made only minor adjustments to the observed dipoles and revealed no significant evidence of systematic effects, it is possible that the BOSS data suffer from unmodeled systematics in spectroscopic redshift determination not present in the eBOSS data. Particularly, adopting the prior $|\beta|c\leq 3200$ km/s, it was not possible to effectively minimize our dipole estimator in the redshift range $0.40.2<z<$ 0.4, corresponding to LOWZ galaxies, although the minimization worked as expected with the BOSS mocks. This discrepancy suggests that our BOSS measurement may be biased. Consequently, we will consider the eBOSS measurement as our principal result.
[…]
  1. 1.
    Our eBOSS redshift dipole measurement is accurate, the anomalous number count estimates are biased by systematics, and the cosmological principle holds;
  2. 2.
    Our dipole measurement is biased and the anomalous number count estimates are accurate, suggesting the presence of anomalously large-scale inhomogeneities;
  3. 3.
    Both measurements are accurate, indicating the presence of exotic inhomogeneities that generate density but not velocity large-scale perturbations.
Only future data will clarify this issue.
Figure 10: Comparison of dipole measurements in Cartesian coordinates with previous estimates.
Table 2: Significance of tension of previous estimations against our results.
Dipole $\boldsymbol{\Delta}_{1}$ Significance of tension with
Measurement $|\boldsymbol{v}|$ (km/s) $l(^{\circ})$ $b(^{\circ})$ eBOSS BOSS
@@
Progress in this area will require careful consideration of potential overlooked systematic factors in analyses based on number counts and redshifts, such as redshift evolution modeling, biases in estimators, or underestimation of uncertainties [46, 47, 48, 49, 50, 51]. Furthermore, validating the analysis pipeline using realistic mock catalogs and conducting blind data analyses, as undertaken in this study, remains critical. The incoming DESI final release data could reduce the errors by a factor of $\sim 5$, by increasing the number of tracers from $1.4$ million to $30$ millions [52]. This enhancement is expected to significantly improve the signal-to-noise ratio and provide a definitive examination of the reference frame throughout cosmic history. However, this study has revealed that subtle redshift systematics, which may be negligible for BAO analyses, can significantly affect the measurement of the dipolar modulation in redshift caused by our relative velocity. Therefore, a robust determination of this effect will require meticulous characterization of the redshift errors in spectroscopic surveys like DESI.
Acknowledgments
It is a pleasure to thank Nathan Secrest, Subir Sarkar, Tomáš Šoltinský, Christophe Yèche, Pierre Fleury and Julián Bautista for useful comments and discussions. Funding: PSF thanks FAPES (Brazil) for financial support. VM thanks CNPq (Brazil) and FAPES (Brazil) for partial financial support. Sci-Com cluster: We acknowledge the use of the computational resources provided by the Sci-Com Lab of the Department of Physics at UFES, which was funded by FAPES. Joint CHE / Milliways cluster: We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. HOTCAT cluster: We acknowledge the use of the HOTCAT computing infrastructure of the Astronomical Observatory of Trieste of the National Institute for Astrophysics (INAF, Italy). MultiDark-Patchy mocks: The massive production of all MultiDark-Patchy mocks for the BOSS Final Data Release has been performed at the BSC Marenostrum supercomputer, the Hydra cluster at the Instituto de Fısica Teorica UAM/CSIC, and NERSC at the Lawrence Berkeley National Laboratory. We acknowledge support from the Spanish MICINNs Consolider-Ingenio 2010 Programme under grant MultiDark CSD2009-00064, MINECO Centro de Excelencia Severo Ochoa Programme under grant SEV- 2012-0249, and grant AYA2014-60641-C2-1-P. The MultiDark-Patchy mocks was an effort led from the IFT UAM-CSIC by F. Prada’s group (C.-H. Chuang, S. Rodriguez-Torres and C. Scoccola) in collaboration with C. Zhao (Tsinghua U.), F.-S. Kitaura (AIP), A. Klypin (NMSU), G. Yepes (UAM), and the BOSS galaxy clustering working group. EZmocks: We acknowledge the use of the EZmocks galaxy mock catalogues with redshift evolution and systematics for galaxies and quasars of the final SDSS-IV data release [35]. SDSS-III (BOSS) data: Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III web site is http://www.sdss3.org/. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University. SDSS-IV (eBOSS) data: Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS web site is www.sdss4.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.